# T-FIX: Text-Based Explanations with Features Interpretable to eXperts

### v1

@@ Abstract
As LLMs are deployed in knowledge‑intensive settings (e.g., surgery, astronomy, therapy), users expect not just answers, but also meaningful explanations for those answers. In these settings, users are often domain experts (e.g., doctors, astrophysicists, psychologists) who require explanations that reflect expert-level reasoning. However, current evaluation schemes primarily emphasize plausibility or internal faithfulness of the explanation, which fail to capture whether the content of the explanation truly aligns with expert intuition. We formalize expert alignment as a criterion for evaluating explanations with T‑FIX, a benchmark spanning seven knowledge-intensive domains. In collaboration with domain experts, we develop novel metrics to measure the alignment of LLM explanations with expert judgment. [^1] https://github.com/BrachioLab/FIX-2
T-FIX: Text-Based Explanations with Features Interpretable to eXperts
[…]
@@ 1 Introduction
Figure 1: Current evaluations of LLM explanations typically consider two dimensions: plausibility — whether the reasoning is logically coherent and faithfulness — whether it reflects the model’s true decision process. We introduce a third, orthogonal dimension: expert alignment — whether the LLM reasons as a domain expert would. For instance, an LLM may correctly predict sepsis risk with a plausible and faithful explanation, yet because it relies on features clinicians rarely use, its expert alignment is low.
@@
LLMs are increasingly used for domain-specific tasks that require substantial knowledge from specialized fields — the adoption of LLMs is actively being explored in high-stakes environments like operating rooms, astronomical observatories, and therapy clinics pressman2024clinical; fouesneau2024rolelargelanguagemodels; stade2024large. For LLMs to be trustworthy in these critical applications, users require not only correct answers but also good explanations rudin2019stop; pedreschi2019meaningful.
However, what constitutes a “good explanation” depends on the target audience ribera2019can; sokol2020one. As LLMs are used for specialized tasks like surgical assistance or supernova analysis, the primary users are often domain experts, like doctors or astrophysicists. Consequently, a “good explanation” in these specialized contexts must offer insights that are valuable and interpretable to these domain experts.
@@
Existing evaluations of LLM explanations focus primarily on two dimensions: (1) plausibility, which assesses whether an explanation appears intuitively reasonable to an average human observer; and (2) faithfulness, which measures whether the explanation accurately reflects the LLM’s internal reasoning process zhou2021evaluating; agarwal2024faithfulness; parcalabescu2023measuring.
@@
While these dimensions are necessary, they are insufficient for knowledge-intensive domains. Experts do not merely ask whether an explanation seems plausible or mirrors the model’s internal reasoning; they require that the reasoning itself aligns with domain-validated principles wang2021explanations. In such settings, the key question is whether LLMs consider aspects of the input that experts themselves deem scientifically critical.
This limitation is illustrated in Figure 1: an LLM may generate a correct answer with a plausible and faithful explanation, yet still rely on features irrelevant to the expert’s reasoning framework. Such reasoning, though logically coherent, fails to meet expert standards. To address this, we introduce a third evaluation dimension: expert alignment.
Expert alignment assesses whether an explanation emphasizes the criteria a domain expert would use when making the same prediction. It is orthogonal to plausibility and faithfulness: while plausibility evaluates logical coherence and faithfulness evaluates causal accuracy, expert alignment tests for epistemic validity — that the explanation is grounded in established scientific knowledge.
While prior work has examined expert reasoning via structured feature attribution jin2024fix, such methods are largely confined to classification settings. In contrast, modern LLMs produce free-form textual explanations not directly linked to explicit feature groups. To our knowledge, no existing benchmark evaluates the alignment of such free-form explanations.
@@
To fill this gap, we introduce the T-FIX benchmark: a suite of datasets spanning seven domains with an accompanying evaluation framework. Developed with domain experts, T-FIX quantifies the extent to which LLM explanations align with scientifically-grounded expert reasoning.
Our contributions are as follows:
@@
  • We introduce expert alignment as a desired attribute of LLM-generated explanations and create T-FIX, the first benchmark designed to evaluate this.
@@
  • We release a pipeline to evaluate how well any LLM “thinks like an expert,” designed to be easily extendable to new domains.
@@
  • We show that while LLMs perform better when reasoning over multiple expert criteria, they still struggle to align their explanations with expert intuition, revealing a key direction for future improvement.
2 Expert Alignment Criteria
@@
Figure 3: Our T-FIX pipeline. To evaluate an LLM-generated explanation, we first decompose it into atomic claims. Next, we filter out irrelevant claims, such as unsupported or speculative statements. Each remaining claim is then scored against the domain-specific expert alignment criteria: a score of “complete” indicates perfect overlap with at least one criterion, while “none” indicates no overlap. Filtered-out claims are automatically assigned a score of “none”. We compute the final expert-alignment score for the explanation by averaging across all claim scores.
@@
When answering knowledge-intensive questions like “Will this patient develop sepsis in the next 12 hours?” or “What kind of supernova produced these wavelengths?”, doctors and astrophysicists rely on domain-specific heuristics, prioritizing certain features over others based on training and experience. For instance, in sepsis classification, an experienced clinician would typically emphasize features like advanced age and hypotension, while assigning lower importance to signals like glucose levels or patient demeanor, which are less directly indicative of sepsis risk.
Thus, an LLM that makes the correct prediction by attending to age and hypotension is more expert-aligned than one that arrives at the same answer by focusing on glucose and demeanor. We define the subset of features that experts prioritize when performing a task as expert alignment criteria.
Step 1: Surveying the field.
To seed our initial list of expert criteria, we prompt OpenAI’s o3 model to perform a comprehensive literature review of the relevant field. Each prompt includes a task description, example input-output pairs from the dataset, and instructions to generate a list of criteria considered important for performing the task – accompanied by reputable citations.
We begin with this deep research approach to avoid over-reliance on any single expert’s perspective. Our goal is to synthesize insights from a broad array of books, journals, and academic publications to produce as comprehensive a list as possible.
Step 2: Iteration with domain experts.
To validate and improve the output from Step 1, we present the preliminary criteria list to a domain expert (see Figure 4 for details on each expert per domain). We ask the expert to (1) remove any incorrect or irrelevant criteria, (2) add any important ones that were missed, and (3) ensure that the list reflects a consensus that their peers would accept. The expert then refines the list until it accurately captures the field’s knowledge.
An example criterion for sepsis classification is as follows: Advanced age (over 65 years) markedly increases susceptibility to rapid sepsis progression and higher mortality after infection.
All Deep Research prompt templates and final expert alignment criteria lists for all domains are available in our GitHub repository.
@@ 3 T-FIX Pipeline
Figure 4: Overview of datasets and domains in T-FIX. We evaluate LLM expert alignment across seven diverse domains, spanning cosmology, psychology, and medicine. For each dataset, we highlight the motivating task, input–output format, representative example, and the expert responsible for validating alignment criteria. The final row summarizes the expert alignment criteria used for scoring explanations in each domain. The column colors reflect dataset modality: blue indicates vision, yellow indicates language, and pink indicates time-series.
LLM-generated explanations contain a mix of reasoning steps – some aligned with expert judgment, and others based on irrelevant information.
To systematically evaluate such complex explanations, we first break them down into atomic claims, or standalone “features” that can be individually assessed for expert alignment. By scoring each feature separately and then aggregating these scores, we can compute an overall expert alignment score for the full explanation. See Figure 3 for an example of this multi-step process.
Our T-FIX pipeline, built using GPT-4o, for evaluating expert alignment consists of three components:
@@
    Relevancy Filtering: Removing claims that are unsupported, speculative, or otherwise irrelevant to the model’s prediction.
@@
    Alignment Scoring: Measuring the degree of overlap between each remaining claim and domain expert criteria.
@@ 3.1 Stage 1: Atomic Claim Extraction
Given a free-form text explanation accompanying an LLM’s prediction, our first goal is to identify and extract the distinct reasoning steps, i.e. “features”, used by the LLM. We achieve this by decomposing the explanation into atomic claims.
An atomic claim is defined as a self-contained, indivisible statement that conveys a single verifiable fact, and can be fully understood without reference to the surrounding context.
To extract atomic claims, we adapt prompting techniques from the claim decomposition literature wanner2024closer; gunjal2024molecularfactsdesideratadecontextualization and prompt GPT-4o to transform a free-form explanation into a list of fully decontextualized atomic claims. We treat each claim as representing a single “feature” in the LLM’s explanation.
3.2 Stage 2: Relevancy Filtering
Not all extracted claims contribute meaningfully to expert reasoning. Some may be unsupported (i.e., references to content not present in the input), speculative (i.e., unfounded hypotheses), or otherwise irrelevant (e.g., repeating the model’s final prediction or citing unrelated information).
Given domain experts’ preference for succinct, informative explanations, we prompt GPT-4o to remove such noisy claims by evaluating each atomic claim based on the original input. A claim is retained if it satisfies the following two criteria:
(1) Clearly grounded in and supported by the input (i.e., not unfounded or speculative)
(2) Directly contributes to explaining why the model made its prediction.
On average, 72% of the claims generated in Stage 1 pass this relevancy filter and are carried forward for alignment scoring.
@@ 3.3 Stage 3: Alignment Scoring
In the final stage of our pipeline, we evaluate each retained atomic claim by comparing it to the domain-specific expert alignment criteria (see Section 2). This step quantifies how the reasoning in the LLM’s explanation reflects expert judgment.
Given an atomic claim and a list of expert criteria, we prompt GPT-4o to measure the claim’s expert alignment in two steps:
  1. 1.
    Identify the most aligned expert criterion. The model selects the criterion whose focus and intent best match the core idea of the atomic claim. The model may also indicate that no criteria align with the claim.
  2. 2.
    Label claim alignment. The model scores how well the claim aligns with the chosen criterion: complete, partial, or none. See Table˜1 for details.
Label Meaning
none The claim references something unrelated to the expert category, or misinterprets the category’s meaning.
partial The claim partially refers to the expert category but lacks key details, uses vague language, is overly general, or contains noise.
complete The claim is specific, directly relevant, and fully captures the meaning and intent of the expert category.
For example, consider the expert criterion for sepsis classification: Advanced age (over 65 years). The claim ‘‘The patient is at risk as they are 72 years old’’ would be labeled complete, as it directly and fully supports the criterion. In contrast, the claim ‘‘The patient is at risk as they are 37’’ may be labeled partial: while it discusses patient age, the specific value does not align with the expert threshold for elevated risk. In contrast, the claim ‘‘The patient is NOT at risk as they are 37’’ would be labeled complete.
Examples of claims with high and low alignment for each domain, along with rationale for why those scores were assigned, are provided in Table˜A4.
3.4 Final Aggregation
We assign an alignment label of none to the claims that were filtered out or did not align with any criteria. This ensures LLM-generated explanations are penalized for unsupported or speculative statements, irrelevant information, and misaligned reasoning. Then, we map alignment labels to numerical scores as follows: complete $\rightarrow 1$, partial $\rightarrow 0.5$, and none $\rightarrow 0$. We average these scores across all claims to produce a final expert alignment score for the explanation.
The prompts for all three stages can be found in Appendix˜B and in our Github repository.
@@ 4 Pipeline Validation
Given our pipeline relies on multiple curated GPT-4o prompts, we want to ensure that the extracted and filtered claims are accurate, and that the final alignment scores match domain expert intuition. To do this, we conduct a quantitative evaluation via annotation, and a qualitative evaluation via domain expert review.
Annotation study.
To quantitatively validate the outputs at each stage, we manually annotate 35 examples (5 per domain), covering 295 extracted claims and 211 aligned claims. Six annotators participate, with two annotators per example. [^2] Annotators are PhD students in machine learning at an American university with experience evaluating LLM outputs.
Annotators evaluate each stage using a unified labeling scheme: (A) fully accurate, (B) partially accurate, or (C) incorrect. For atomic claim extraction, they review the explanation and its extracted claims, labeling outputs as (A) correct decomposition, (B) 1–2 claims missing/incorrect, or (C) 3+ claims missing/incorrect. For relevancy filtering, they assess whether claims were appropriately kept or removed, labeling results as (A) all correct, (B) 1–2 incorrect, or (C) 3+ incorrect. For expert alignment scoring, annotators review the filtered claims and expert categories, labeling each as (A) matches expert judgment, (B) borderline, or (C) does not match expert judgment.
These labels are then mapped to accuracy scores of 1.0, 0.5, and 0.0, respectively. Table˜2 reports average accuracy and Cohen’s $\kappa$ across all seven T-FIX domains, indicating moderate-to-substantial agreement and supporting the reliability of our pipeline. Domain-level metrics appear in Table˜A3.
Domain expert interview.
To qualitatively validate the expert alignment scores in Stage 3 of our pipeline, we interview each domain expert and have them analyze 10-20 claims for their dataset, discussing noticed successes and failures.
Across domains, experts largely agreed with the LLM alignment labels, confirming that the evaluation framework reliably captures expert reasoning. Agreement was highest in domains with well-defined interpretive criteria (e.g., physics and psychology) and somewhat lower in medicine, where decomposing multi-symptom medical statements occasionally caused partial rather than full alignment. Overall, expert validation supports the interpretability and trustworthiness of the T-FIX evaluation pipeline. We provide a summary of each domain expert’s interview in Appendix˜E.
Pipeline Stage $\mathcal{N}$ Accuracy Cohen’s $\kappa$
Claim Extraction 35 0.943 0.717
Relevancy Filtering 295 0.871 0.402
Expert Alignment 211 0.923 0.405
Table 2: Pipeline validation: Average accuracy for $\mathcal{N}$ samples across all T-FIX domains and annotator agreement – Cohen’s $\kappa$ for each stage in our pipeline. Domain-specific statistics are provided in Table˜A3.

### v2

@@ Abstract
As LLMs are deployed in knowledge-intensive settings (e.g., surgery, astronomy, therapy), users are often domain experts who expect not just answers, but explanations that mirror professional reasoning. However, most automatic evaluations of explanations prioritize plausibility or faithfulness, rather than testing whether an LLM “thinks like an expert.” Existing approaches to evaluating professional reasoning rely heavily on per-example expert annotation, making such evaluations costly and difficult to scale. To address this gap, we introduce the T-FIX benchmark, spanning seven scientific tasks across three domains, to operationalize expert alignment as a desired attribute of LLM-generation explanations. Our framework enables automatic evaluation of expert alignment, generalizing to unseen explanations and eliminating the need for ongoing expert involvement. Code and data available at https://github.com/BrachioLab/FIX-2
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Current automatic evaluations of LLM explanations focus on plausibility — whether the reasoning is logically coherent and faithfulness — whether it reflects the model’s true decision process. We formalize a third, orthogonal dimension: expert alignment — whether the LLM reasons as a domain expert would. For instance, an LLM may correctly predict sepsis risk, yet because it relies on features clinicians rarely use, its expert alignment is low.
@@
LLMs are increasingly used for domain-specific tasks that require substantial knowledge from specialized fields — the adoption of LLMs is actively being explored in high-stakes environments like operating rooms, astronomical observatories, and therapy clinics (Pressman et al., 2024; Fouesneau et al., 2024; Stade et al., 2024). For LLMs to be trustworthy in these critical applications, users require not only correct answers but also good explanations (Rudin, 2019; Pedreschi et al., 2019).
However, what constitutes a “good explanation” depends on the target audience (Ribera and Lapedriza, 2019; Sokol and Flach, 2020). In specialized tasks such as surgical assistance or supernova analysis, the primary users are domain experts (e.g., doctors or astrophysicists). “Thus, a good explanation” must surface the scientifically relevant criteria experts use in practice. This paper introduces T-FIX to measure this at scale: a suite of datasets and an expert-designed evaluation framework that automatically measures how much an explanation aligns with expert intuition and generalizes to unseen examples across scientific domains.
@@
Current automatic evaluations of LLM explanations typically focus on two dimensions: (1) plausibility, whether an explanation appears intuitively reasonable to an average observer; and (2) faithfulness, whether it reflects the LLM’s internal reasoning process (Zhou et al., 2021; Agarwal et al., 2024; Parcalabescu and Frank, 2023).
@@
While necessary, these dimensions are insufficient in knowledge-intensive domains, where experts require reasoning that aligns with domain-validated principles (Wang and Yin, 2021); the key question is whether LLMs attend to aspects of the input that experts deem scientifically critical. As Figure 1 illustrates, an LLM may produce a correct answer with a plausible and faithful explanation, yet still rely on features irrelevant to expert reasoning; such explanations are coherent but fail to meet expert standards.
We argue for automatic evaluation of a third dimension, expert alignment, which measures whether an explanation emphasizes the criteria a domain expert would use, and tests epistemic validity — grounding in established scientific knowledge, rather than coherence or causal fidelity.
Current methods for measuring expert alignment of LLM outputs across medicine, psychology, and astronomy are not automatic or scalable because they require substantial manual labor from the expert. For example, Kutbi et al. (2025) annotate only 18 sepsis-diagnosis explanations with 7 physicians, and Hyk et al. (2025) report four weeks of interviews with 11 astronomers. These studies are difficult to turn into benchmarks: the evaluations are limited in size and often qualitative, and therefore do not generalize beyond the specific annotated examples.
Conversely, larger benchmarks for domain-specific LLM evaluation rarely assess expert alignment in a substantive way. In many cases, experts appear primarily in validation or annotation studies (e.g., Astro-QA (Li et al., 2025a), CounselBench (Li et al., 2025b), Mentalchat16k (Xu et al., 2025)). In medicine, newer efforts introduce per-example, domain-specific rubrics (e.g., Health-SCORE from Yang et al. (2026) and OpenAI’s HealthBench), but they share a key limitation: expert input is tied to the benchmark’s fixed examples and does not readily transfer to evaluating explanations on unseen cases.
@@
To fill this gap, we introduce the T-FIX benchmark: a suite of seven datasets spanning three scientific domains with an accompanying expert-designed evaluation framework. Developed with domain experts, T-FIX quantifies how well LLM explanations align with scientifically grounded expert reasoning. Our contributions are as follows:
@@
  • We formalize expert alignment as a desired attribute of LLM-generated explanations and create T-FIX, the first benchmark designed to evaluate this automatically.
@@
  • We release an expert-designed framework to evaluate how well any LLM “thinks like an expert,” across seven unique scientific tasks.
@@
  • We show that while LLMs perform better when reasoning over multiple expert criteria, they do not explicitly rely on expert reasoning, revealing a key direction for future improvement.
T-FIX & Scalability. With T-FIX, expert effort is front-loaded and reduced: experts help develop a single, robust evaluation framework, after which users can score expert alignment on unseen LLM explanations automatically, eliminating the need for future per-example annotations or additional expert time.
2 Expert Criteria for Evaluating Explanations
@@
Figure 3: Our T-FIX pipeline. To evaluate an LLM-generated explanation, we first decompose it into atomic claims. Next, we filter out irrelevant claims (e.g., unsupported or speculative statements) and group the remaining claims by their relation to each expert criterion. We then score each criterion-level group against its corresponding expert criterion as complete, partial, or none. Each claim inherits the highest score among the groups it belongs to, while filtered-out or ungrouped claims are assigned none. We compute the final expert alignment score by averaging these per-claim scores across the explanation.
@@
When answering knowledge-intensive questions like “Will this patient develop sepsis in the next 12 hours?” or “What kind of supernova produced these wavelengths?”, doctors and astrophysicists rely on domain-specific heuristics, prioritizing certain features over others based on their training and experience.
For instance, in sepsis classification, an experienced clinician would typically emphasize features like advanced age and elevated temperature, while assigning lower importance to signals like patient demeanor, which are less directly indicative of sepsis risk.
Thus, an LLM that makes the correct prediction using age and temperature is more expert-aligned than one that arrives at the same answer by focusing on only patient demeanor. We define the features that experts prioritize as expert alignment criteria. We draw inspiration from prior work (Yang et al., 2026; Li et al., 2025b), but seek to move beyond example-level rubrics to task-level criteria.
Developing Expert Criteria. We work with five medical experts (surgeons and clinicians), three cosmologists (astrophysicists and astronomers), and two psychologists to develop the expert criteria for each of our datasets. See Figure 4 for details on collaborating experts.
To save our domain experts’ time, we seed our initial list of expert criteria using o3 deep research (OpenAI, 2025c) to perform a comprehensive literature review of the relevant field. Each prompt includes a task description, examples, and instructions to generate a list of criteria considered important for performing the task – accompanied by reputable citations from academic books, journals, and conferences.
Then, we present the preliminary criteria list to the relevant domain experts and ask them to (1) remove any incorrect or irrelevant criteria, (2) add any important ones that were missed, and (3) ensure that the list reflects a consensus among them. The experts then iteratively refine the list until they feel it accurately captures the field’s knowledge.
An example criterion for sepsis classification is as follows: Advanced age (over 65 years) increases susceptibility to rapid sepsis progression and higher mortality after infection.
All Deep Research prompt templates and final expert alignment criteria are available in our GitHub repository.
@@ 3 T-FIX Pipeline
Next, we develop a framework to use these expert criteria to automatically evaluate explanations. LLM-generated explanations typically contain a mix of reasoning steps — some aligned with expert judgment, and others based on irrelevant information.
To systematically evaluate such complex explanations, we first break them down into atomic claims, or standalone “features.” We then group these claims by expert criteria, score each feature group for expert alignment, and aggregate these scores to compute an overall alignment score for the full explanation. See Figure 3 for an example of this multi-step evaluation process.
Our T-FIX pipeline for evaluating expert alignment consists of three components:
@@
    Claim Filtration & Grouping: Removing claims that are irrelevant to the model’s prediction and grouping claims by relation to each expert criterion.
@@
    Alignment Scoring: Measuring how much each group of claims supports its corresponding expert criteria.
We implement this pipeline using an evaluator LLM; results reported below use GPT-5-mini (OpenAI, 2025b).
@@ 3.1 Stage 1: Atomic Claim Extraction
Given a free-form text explanation accompanying an LLM’s prediction, our first goal is to identify and extract the distinct reasoning steps, i.e. “features”, used by the LLM. We achieve this by decomposing the explanation into atomic claims. An atomic claim is defined as a self-contained, indivisible statement that conveys a single fact, and can be fully understood without reference to the surrounding context.
To extract atomic claims, we adapt prompting techniques from the claim decomposition literature (Wanner et al., 2024; Gunjal and Durrett, 2024) and prompt GPT-5-mini to transform a free-form explanation into a list of fully decontextualized atomic claims. We treat each claim as representing a single “feature” in the LLM’s explanation.
3.2 Stage 2: Claim Filtering & Grouping
Filtering. Not all extracted claims contribute meaningfully to expert reasoning. Some may be unsupported (i.e., references to content not in the input), speculative (i.e., unfounded hypotheses), or otherwise irrelevant (e.g., repeating the model’s final prediction or citing unrelated information).
Given domain experts’ preference for succinct, informative explanations, we prompt GPT-5-mini to remove such noisy claims by evaluating each atomic claim based on the original input. A claim is retained if it is (1) Clearly grounded in and supported by the input, and (2) Directly contributes to explaining why the model made its prediction. On average, 76.8% of the claims generated in Stage 1 pass this relevancy filter and are carried forward for grouping.
Grouping. Next, we prompt GPT-5-mini to group extracted claims by their relationship to each expert criterion. For each criterion, the LLM selects the subset of claims that are topically related to that criterion.
Formally, given a set of expert criteria $K=\{k_{1},\ldots,k_{m}\}$ and an explanation consisting of claims $C=\{c_{1},\ldots,c_{n}\}$, we define the claim group of criterion $k$ as:
$G_{k}=\{c\in C:\mathsf{related}(c,k)\},$ (1)
where we use GPT-5-mini to check the $\mathsf{related}$ predicate.
Importance of Claim Grouping. We group claims at the criterion level because expert alignment is often determined by the joint content of multiple claims: claims can depend on one another, and some criteria are inherently compositional. For example, SIRS positivity in sepsis diagnosis requires meeting at least two symptoms from a predefined list, which may be expressed across several separate claims.
Label Meaning
complete The claims are all specific, directly relevant, and fully capture the meaning and intent of the expert criterion.
partial The claims partially support the expert criterion but lack key details, use vague language, are overly general, or contain noise.
none The claims reference something unrelated to the expert criterion, misinterprets the criterion’s meaning, or contradict what is stated in the criterion.
@@ 3.3 Stage 3: Alignment Scoring
In the final stage of our pipeline, we score each criterion-level group of claims by assessing how well the group, taken as a whole, reflects its corresponding domain-specific expert criterion. That is, we wish to assign a numerical score to each claim group $G_{k}$. We compute this score by prompting the LLM GPT-5-mini to label the group’s alignment as complete, partial, or none (see Table˜1), and then assigning these labels to numerical values.
For example, consider the expert criterion for sepsis classification: Advanced age (over 65 years). The claim ‘‘The patient is at risk as they are 72 years old’’ would be labeled complete, as it directly and fully supports the criterion. The claim ‘‘The patient is NOT at risk as they are 37’’ would also be labeled complete. In contrast, the claim ‘‘The patient is at risk as they are 64’’ may be labeled partial: It reasons about patient age correctly, but the specific value does not align with the expert threshold for elevated risk.
We express the LLM-based labeling of a claim group with respect to its expert criterion, $\mathcal{L}(G_{k},k)\in\{\textit{complete},\textit{partial},\textit{none}\}$, as the group-wise alignment:
[…]
Examples of claims with high and low alignment for each domain, along with rationale for why those scores were assigned, are provided in Table˜A5.
3.4 Final Aggregation: T-FIXscore
Now that we’ve measured the alignment of each claim group, we turn to the harder task of isolating the alignment score of each individual claim.
We calculate claim-level alignment scores by taking the maximum score across all groups that contain the claim. Any claim that supports at least one criterion is thus treated as aligned with expert intuition.
For the set of claims $C$, criteria $K$, and the corresponding criterion-level claim groups $\{G_{1},\dots,G_{m}\}$, we formally define claim alignment as:
$\mathsf{CAlign}(c_{i},K)=\max_{k\in K:c_{i}\in G_{k}}\big(\mathsf{GAlign}(G_{k},k)\big)$ (3)
where we let the maximum over the empty set be 0, such as in the case where a claim is unrelated to any criterion. This penalizes irrelevant information and misaligned reasoning.
Finally, we average the per-claim alignment score to compute the final explanation alignment score as:
$\mathsf{T\textrm{-}FIXscore}(C,K)=\frac{1}{n}\sum_{i=1}^{n}\mathsf{CAlign}(c_{i},K)$ (4)
We draw inspiration from Jin et al. (2025), which details a method for aggregating expert alignment for input feature groups typically extracted from trained classification models. With $\mathsf{T\textrm{-}FIXscore}$, we extend the metric that Jin et al. (2025) proposes to free-form text explanations from generative models not directly linked to explicit feature groups.
3.5 Prompt Development with Expert Feedback
We met with our domain experts throughout the prompt development process for each step of our pipeline. We asked what they would do differently from the LLM, incorporated their feedback, and iterated on all prompts multiple times, until the experts were satisfied with the results.
Our prompt iteration process is detailed in Appendix F. The final prompts for all three stages can be found in Appendix˜E and in our Github repository.
@@ 4 Pipeline Validation
Given our pipeline relies on multiple curated GPT-5-mini prompts, we want to ensure that the extracted and filtered claims are accurate, and that the final alignment scores match domain expert intuition. To do this, we conduct a quantitative evaluation via an annotation study, and a qualitative evaluation via domain expert interviews.
Annotation study. To quantitatively validate the outputs at each stage, we manually annotate 35 examples (5 per domain), covering 295 extracted claims and 211 aligned claims. Six annotators participate, with two annotators per example. [^1] Annotators are PhD students in machine learning trained extensively by relevant domain experts to evaluate LLM outputs. See Appendix B for details on study setup. Table˜A1 reports average accuracy and Cohen’s $\kappa$ across all seven T-FIX datasets, indicating substantial agreement and supporting the reliability of our pipeline. Domain-level metrics appear in Table˜A3.
Domain expert interviews. To qualitatively validate the expert alignment scores in Stage 3 of our pipeline, we conduct ongoing interviews with each domain expert and have them annotate and analyze 10-20 claims for their dataset, discussing noticed successes and failures.
Across domains, experts largely agreed with the LLM alignment labels, confirming that the evaluation framework reliably captures expert reasoning. Agreement was highest in domains with well-defined interpretive criteria (e.g., physics and psychology). Additional prompt iteration incorporating interview feedback was required to achieve high agreement in medicine, as multi-symptom, compositional medical statements were more complex. Overall, expert validation supports the interpretability and trustworthiness of the T-FIX evaluation pipeline. We provide a summary of each domain expert’s interview in Appendix˜D.
Figure 4: Overview of datasets and domains in T-FIX. We evaluate LLM expert alignment across seven diverse scientific tasks, spanning cosmology, psychology, and medicine. For each dataset, we highlight the motivation, I/O format, representative example, and the key expert responsible for validation. The final row summarizes the expert alignment criteria used for scoring explanations in each domain. The column colors reflect dataset modality: blue indicates vision, yellow indicates language, and pink indicates time-series.